# VaR at Its Extremes: Impossibilities and Conditions for One-Sided Random Variables

### v2

@@ 1 Introduction
The study of quantiles has long been a cornerstone of mathematical and statistical research. Quantiles provide a fundamental link between abstract probability models and the observable outcomes they generate. In risk management, in particular, quantiles have been regarded as essential tools for assessing the riskiness of losses, asset prices, and other financial variables. Their most prominent manifestation is the Value-at-Risk (VaR) [Linsmeier2000], which is interpreted as the minimum amount of capital a financial institution must hold so that losses exceed this level only with a small, pre-specified probability.
A variety of risk measures have been developed to refine or extend VaR. Among them, the conditional tail expectation – also known as expected shortfall – [Acerbi.2002, Tasche2002] which addresses several limitations of VaR, most notably the failure of sub-additivity. Nevertheless, interest in VaR has persisted, partly due to its ability to capture the opposite phenomenon of super-additivity, a feature that does not rely on the integrability requirements imposed by alternative risk measures. In this paper, we examine the extremal behaviors of VaR, focusing on the conditions under which it exhibits sub-additivity or super-additivity across all probability thresholds.
To formalize our analysis, we consider random vectors $\bm{X}=(X_{1},\dots,X_{n})$, $n\in\mathbb{N}$, whose components are random variables representing, for example, asset prices or insurance losses. We focus in particular on their aggregate,
@@
In the final section, Section 4, we show how this assumption can be relaxed. In particular, we extend all results to the setting where the lower endpoints $a_{i}$ are arbitrary but finite, and also to the complementary case in which the random variables are instead bounded above.
@@
Our primary objective is to investigate how the VaR of the sum, ${\mathrm{VaR}}_{p}[S]$, compares to the sum of the individual VaRs, $\sum_{i=1}^{n}{\mathrm{VaR}}_{p}[X_{i}]$ for all $p\in(0,1)$. To provide a precise framework for this comparison, we introduce the following definitions.
@@
The remainder of the paper is organized as follows. Section 2 establishes an impossibility theorem for VaR sub-additivity, extending the recent findings of [Imamura2025] and showing that sub-additivity can occur only in the degenerate case of VaR additivity. Section 3 develops a new and unified characterization of VaR super-additivity that encompasses most existing results in the literature while allowing for non-identically distributed margins and a wider range of dependence structures. In Section 4, we generalize these results to random variables with arbitrary finite lower or upper endpoints. Section 5 concludes the paper.
@@ 2 VaR Sub-additivity
VaR sub-additivity is widely regarded as a desirable property, as it reflects the risk-reducing effect of diversification. In pursuit of this property, numerous alternative risk measures have been introduced to guarantee it. The literature has examined VaR sub-additivity in various settings, including asymptotic regimes [Danielsson2013] and classes of distributions such as the elliptical distributions where VaR is known to be sub-additive for confidence levels $p\geq\frac{\displaystyle 1}{\displaystyle 2}$ [McNeil2015]. Nevertheless, when we examine VaR sub-additivity for right-sided random variables, the conclusion turns out to be remarkably simple. As we show in the next theorem, VaR sub-additivity cannot occur in our setting except in the degenerate case of exact additivity. Before presenting this main result, we recall the notion of co-monotonicity, which represents the extremal form of positive dependence.
@@
$S_{k}\stackrel{{\scriptstyle d}}{{\coloneqq}}\sum_{i=1}^{n}X_{i,k}.$
The conditional variables $X_{i,k}$ are well-defined since each $X_{i}$ has a zero lower endpoint and $k>0$.
@@
For $x_{i}<{\mathrm{VaR}}_{F_{S}(k)}[X_{i}]$, we have $F_{X_{i}}(x_{i})<F_{S}(k)\leq F_{X_{i}}({\mathrm{VaR}}_{F_{S}(k)}[X_{i}])$. Then the ratio $\frac{\displaystyle F_{X_{i}}(x_{i})}{\displaystyle F_{S}(k)}$ is strictly less than 1, and $x_{i}<k$, so the CDFs are well-defined.
@@
Since each $X_{i,k}$ has a finite expectation ($\mathbb{E}[X_{i,k}]\leq k<\infty$), Theorem 1 in [Imamura2025] implies that $\bm{X}_{k}$ is co-monotonic and consequently VaR additive i.e.:
@@
Theorem 2.2 extends and strengthens the main result of [Imamura2025], which relied on integrability assumptions. It reveals the rigid phenomenon of the impossibility of VaR sub-additivity for random variables supported on $0,\infty)$. The sub-additivity inequality is satisfied only in the degenerate case where VaR is exactly additive, and such additivity occurs exclusively under co-monotonicity.
@@ 3 VaR super-additivity
Unlike the sub-additivity property of VaR, the opposite effect -— VaR super-additivity -— can in fact arise. For instance, consider the case of a counter-monotonic random vector.
@@
whose joint CDF is given by
@@
For simplicity, take $\alpha=\theta=1$. The marginals are then identical Pareto(II) variables with
${\mathrm{VaR}}_{p}[X]=\frac{p}{1-p}.$
@@
${\mathrm{VaR}}_{p}[X+1/X]=\frac{2(1+p^{2})}{1-p^{2}}.$
@@
$\frac{2(1+p^{2})}{1-p^{2}}>\frac{2p}{1-p},\qquad\forall\,p\in(0,1),$
@@
Although several studies have constructed families of distributions that exhibit tail-level VaR super-additivity [Embrechts2008, Embrechts2009b, Zhu2023], relatively few have examined this behavior across all probability levels $p$. [Ibragimov2009] analyzed full-range super-additivity in i.i.d. stable distributions, and [Chen2025] introduced a class of random vectors with identically distributed, weakly negatively associated margins for which VaR is fully super-additive. The counter-monotonic vector presented above is a special case of this class. Taken together, these results show that VaR super-additivity is not a pathological anomaly but occurs naturally under economically meaningful and probabilistically coherent conditions.
Although not stated in this form, Theorem 1 of [Imamura2025] also implies a strong constraint on when super-additivity can happen. If all components $X_{i}$ are integrable i.e. $\mathbb{E}[|X_{i}|]<\infty$, then VaR super-additivity is equivalent to VaR additivity. Consequently, to construct examples of random vectors $\bm{X}$ that are genuinely VaR super-additive, at least one of the components must be non-integrable. Our counter-monotonic example above shows this implication as the two Pareto II variables have infinite means. The converse, however, does not necessarily hold. Non-integrability is intrinsically a tail property and does not by itself guarantee full super-additivity. The following example illustrates this point.
@@
    we obtain $\mathbb{E}\!\left[\frac{\displaystyle 1}{\displaystyle\mathstrut 1+X}\right]=\frac{\displaystyle 1}{\displaystyle 2}$, while $\mathbb{E}[X]=\infty$ for this Pareto distribution.
@@
    For the sum $S=X+\frac{\displaystyle 1}{\displaystyle\mathstrut 1+X}$, one can show that
@@
    Comparing ${\mathrm{VaR}}_{p}[S]$ with ${\mathrm{VaR}}_{p}[X]+{\mathrm{VaR}}_{p}\!\left[\frac{\displaystyle 1}{\displaystyle\mathstrut 1+X}\right]$ reveals that VaR is super-additive for $p\in\left(0,\frac{\displaystyle 1}{\displaystyle 2}\right],$ and sub-additive for $p\in\left[\frac{\displaystyle 1}{\displaystyle 2},1\right).$
@@
    where $F_{X_{1}}$ and $F_{X_{2}}$ are Pareto (II) marginal distributions with parameters $\alpha_{1}=\alpha_{2}=\theta_{1}=\theta_{2}=1$. The copula $C(u,v)$ is an Ordinal Sum copula (see Example 3.4 in [10.5555/1952073]) given by
    $C(u,v)=\begin{cases}\max\!\left\{u+v-\frac{\displaystyle 1}{\displaystyle 2},\,0\right\},&(u,v)\in\left[0,\frac{\displaystyle 1}{\displaystyle 2}\right]^{2},\\[5.10011pt] \max\!\left\{u+v-1,\,\frac{\displaystyle 1}{\displaystyle 2}\right\},&(u,v)\in\left(\frac{\displaystyle 1}{\displaystyle 2},1\right]^{2},\\[5.10011pt] \min\{u,v\},&\text{otherwise}.\end{cases}$
@@
    Since both marginals are Pareto (II) with unit shape then their expectations are infinite. Their common VaR is
@@
    ${\mathrm{VaR}}_{p}[S]=\begin{cases}\dfrac{6+8p^{2}}{\,9-4p^{2}\,},&0<p\leq\frac{\displaystyle 1}{\displaystyle 2},\\[6.80011pt] \dfrac{2-2p(1-p)}{p(1-p)},&\frac{\displaystyle 1}{\displaystyle 2}<p<1.\end{cases}$
@@
Example 3.1 demonstrates that even in cases where we have (1) counter-monotonic dependence with one non-integrable margin, and (2) two non-counter-monotonic, non-integrable margins, the resulting dependence-margin combination may still exhibit intervals of VaR sub-additivity. Thus, neither a particular dependence structure nor the mere non-integrability of margins is sufficient on its own to guarantee VaR super-additivity. In fact, the negative dependence used in part (2) of the example is significantly weaker than full counter-monotonicity.
@@
Our objective, therefore, is to identify a dependence property together with a corresponding marginal behavior that, when combined, imply VaR super-additivity. Such a characterization must be sufficiently general to encompass all known results in the literature, yet specific enough to allow for straightforward verification.
@@
is SD then $\bm{X}$ is VaR super-additive.
@@
When $\bm{X}$ is negative lower orthant dependent (NLOD) [Block1982a, Joe1997], that is
@@
The NSD property captures the dependence requirement on the joint distribution of $\bm{X}$ that ensures VaR super-additivity. We note, in passing, that the dependence structure used in part (2) of Example 3.1 is weaker than NSD, specifically it fails the NSD property at $t\in\left(\frac{\displaystyle 7}{\displaystyle\mathstrut 10},1+\sqrt{2}\right)$. This contributed, though was not strictly required, to the failure of VaR super-additivity in that example. Nonetheless, by definition, NSD is a relatively weak form of negative dependence and is strictly implied by NLOD.
@@
${\mathrm{VaR}}_{p}[X]=\frac{p}{1-p},$
@@
We collect below several familiar continuous distributions whose associated functions $\phi_{i}$ are non-increasing.
@@
    $F_{X_{i}}(x_{i})=\exp\left(-\left(\frac{\displaystyle x_{i}}{\displaystyle\theta_{i}}\right)^{-\alpha_{i}}\right),\qquad x_{i}\geq 0,\ \alpha_{i},\theta_{i}>0,$
@@
    $F_{X_{i}}(x_{i})=1-\left(\frac{\displaystyle\theta_{i}}{\displaystyle\theta_{i}+x_{i}}\right)^{\alpha_{i}},\qquad x_{i}\geq 0,\ \alpha_{i},\theta_{i}>0,$
@@
    $\phi_{i}(x_{i})=x_{i}\log\!\left(1-\left(\frac{\displaystyle\theta_{i}}{\displaystyle\theta_{i}+x_{i}}\right)^{\alpha_{i}}\right).$
@@
    $\phi_{i}^{{}^{\prime}}(x_{i})$ $=\log F_{X_{i}}(x_{i})+\frac{\displaystyle x_{i}f_{X_{i}}(x_{i})}{\displaystyle F_{X_{i}}(x_{i})}$
    $=\log F_{X_{i}}(x_{i})+\frac{\displaystyle\alpha_{i}x_{i}\overline{F}_{X_{i}}(x_{i})}{\displaystyle(\theta_{i}+x_{i})F_{X_{i}}(x_{i})}.$
@@
    $\lim_{x_{i}\to\infty}\frac{\displaystyle\phi_{i}^{{}^{\prime}}(x_{i})}{\displaystyle\overline{F}_{X_{i}}(x_{i})}=\alpha_{i}-1.$
@@
    $\phi_{i}^{{}^{\prime}}(x_{i})$ $=-\int_{F_{X_{i}}(x_{i})}^{1}\frac{\displaystyle 1}{\displaystyle w}\,\mathrm{d}w+\frac{\displaystyle\alpha_{i}x_{i}}{\displaystyle(\theta_{i}+x_{i})F_{X_{i}}(x_{i})}\int_{F_{X_{i}}(x_{i})}^{1}\mathrm{d}w$
    $=-\int_{F_{X_{i}}(x_{i})}^{1}\frac{\displaystyle(\theta_{i}+x_{i})F_{X_{i}}(x_{i})-\alpha_{i}x_{i}w}{\displaystyle w(\theta_{i}+x_{i})F_{X_{i}}(x_{i})}\,\mathrm{d}w.$
@@
    Applying the mean value theorem to $t\mapsto t^{\alpha_{i}}$ on $\left[\frac{\displaystyle\theta_{i}}{\displaystyle\theta_{i}+x_{i}},1\right]$ yields the inequality and thus the desired non-positivity.
@@
    $\phi_{i}^{\prime}(x_{i})=\log\!\left(\mathrm{erfc}\!\left(\sqrt{\frac{\theta_{i}}{2x_{i}}}\right)\right)+\frac{\sqrt{\frac{\displaystyle\theta_{i}}{\displaystyle 2\pi x_{i}}}\exp\!\left(-\frac{\displaystyle\theta_{i}}{\displaystyle 2x_{i}}\right)}{\mathrm{erfc}\!\left(\sqrt{\frac{\displaystyle\theta_{i}}{\displaystyle 2x_{i}}}\right)}.$
    Let $t=\sqrt{\frac{\displaystyle\theta_{i}}{\displaystyle 2x_{i}}}$. Then $\phi_{i}^{\prime}(x_{i})\leq 0$ is equivalent to $\psi_{i}(t)\leq 0$, where
@@
    the classical Mills ratio bound [Mills1926]. Thus $\phi_{i}^{\prime}(x_{i})\leq 0$ for all $x_{i}\geq 0$ i.e. $\phi_{i}$ is non-increasing on $0,\infty)$.
@@
    $F_{X_{i}}(x_{i})=\left(\frac{\displaystyle x_{i}}{\displaystyle 1+x_{i}}\right)^{\alpha_{i}},\qquad x_{i}\geq 0,\ \alpha_{i}>0,$
@@
    $\phi_{i}(x_{i})=\alpha_{i}x_{i}\log\!\left(\frac{\displaystyle x_{i}}{\displaystyle 1+x_{i}}\right).$

### v3

@@ 1 Introduction
The study of quantiles has long been a cornerstone of mathematical and statistical research. Quantiles provide a fundamental link between abstract probability models and the observable outcomes they generate. In risk management, in particular, quantiles have been regarded as essential tools for assessing the riskiness of losses, asset prices, and other financial variables. Their most prominent manifestation is the Value-at-Risk (VaR) (Linsmeier and Pearson, , 2000), which is interpreted as the minimum amount of capital a financial institution must hold so that losses exceed this level only with a small, pre-specified probability.
A variety of risk measures have been proposed to refine or extend VaR. Among them, the conditional tail expectation – also known as expected shortfall – (Acerbi and Tasche, , 2002; Tasche, , 2002) which addresses several of its well-known limitations, most notably its potential violation of sub-additivity. Nevertheless, VaR has continued to attract considerable attention in both finance and actuarial science. One reason is its distinctive ability to capture situations in which risk aggregation leads to super-additivity, reflecting the possibility that diversification may fail in the presence of heavy-tailed risks. While this phenomenon lies outside the integrability framework underlying many prominent risk measures, it provides important insights into the behavior of extreme losses that arises naturally in insurance applications (Embrechts et al., , 1997). In this paper, we investigate the extremal aggregation behavior of VaR and identify conditions under which it exhibits sub-additivity or super-additivity uniformly across all probability levels.
The contributions of this paper are threefold. First, we establish an impossibility theorem for VaR sub-additivity (Theorem 2.2). Specifically, for random variables supported on $0,\infty)$, VaR sub-additivity and exact additivity are equivalent, where both can occur exclusively under co-monotonicity. This result extends the main finding of Imamura and Kato, (2025), which was derived under the assumption of finite expectations, to the more general setting that allows for risks with infinite means. Second, we provide a new characterization of full VaR super-additivity (Theorem 3.5) based on two easily verifiable structural conditions: negative simplex dependence imposed on the joint distribution, and simplex dominance imposed on a margin-dependent aggregator function. In contrast to earlier contributions in the literature (see, for example, Chen et al., (2025); Chen and Shneer, (2026); Arab et al., (2025); Müller, (2025)), our framework accommodates non-identically distributed margins and encompasses a broad class of negative dependence structures beyond negative lower orthant dependence. Third, we extend the analysis to random variables with arbitrary finite endpoints, whether lower or upper, and establish an exact duality between the sub-additive and super-additive regimes.
To formalize our analysis, we consider random vectors $\bm{X}=(X_{1},\dots,X_{n})$, $n\in\mathbb{N}$, whose components are non-degenerate random variables representing, for example, asset prices or insurance losses. We focus in particular on their aggregate,
@@
In the final section, Section 4, we show how this assumption can be relaxed. In particular, we extend all results to the setting where the lower endpoints $a_{i}\in\mathbb{R}$ are arbitrary but finite, and also to the reflected setting in which the random variables are instead bounded above.
@@
Our primary objective is to investigate how the VaR of the sum, ${\mathrm{VaR}}_{p}[S]$, compares to the sum of the individual VaRs, $\sum_{i=1}^{n}{\mathrm{VaR}}_{p}[X_{i}]$ for all $p\in(0,1)$. In actuarial practice, such aggregation questions arise naturally when insurers combine heterogeneous lines of business – such as catastrophe, liability, and operational risks – which may depend upon one another and whose marginal distributions may differ substantially. To provide a precise framework for this comparison, we introduce the following definitions.
@@
The remainder of the paper is organized as follows. Section 2 examines VaR sub-additivity for non-negative risks and identifies the conditions under which it arises. Section 3 analyzes VaR super-additivity and develops the framework used to characterize this phenomenon. Section 4 broadens the analysis to random variables with arbitrary finite lower endpoints and to the complementary case of variables bounded above. Section 5 concludes the paper.
@@ 2 VaR Sub-additivity
VaR sub-additivity is widely regarded as a desirable property, as it reflects the risk-reducing effect of diversification. In pursuit of this property, numerous alternative risk measures have been introduced to guarantee it. The literature has examined VaR sub-additivity in various settings, including asymptotic regimes (Daníelsson et al., , 2013) and classes of distributions such as the elliptical distributions where VaR is known to be sub-additive for confidence levels $p\geq\dfrac{1}{2}$ (McNeil et al., , 2015). A recent and simple result of Imamura and Kato, (2025) shows that, for any random variables with finite expectations, imposing VaR sub-additivity at all confidence levels is equivalent to co-monotonicity – the extremal form of positive dependence – and hence to exact VaR additivity. This equivalence still holds, stated in Theorem 2.2 below, for the full class of non-negative random variables, with no moment conditions whatsoever. Our proof constructs a truncated family $\bm{X}_{k}$ of integrable approximations, applies the result of Imamura and Kato, (2025) to each, and recovers the original vector via a monotone convergence argument. Before presenting this result, we recall the notion of co-monotonicity (Dhaene et al., , 2002).
@@
$S_{k}=\sum_{i=1}^{n}X_{i,k}.$
Since each $X_{i}$ has lower endpoint zero, $S\geq 0$ and therefore $\mathbb{P}(S\leq k)>0$ for any $k>0$, so the conditional variables $X_{i,k}$ are well defined.
@@
For $x_{i}<{\mathrm{VaR}}_{F_{S}(k)}[X_{i}]$, we have $F_{X_{i}}(x_{i})<F_{S}(k)\leq F_{X_{i}}({\mathrm{VaR}}_{F_{S}(k)}[X_{i}])$. Then the ratio $\dfrac{F_{X_{i}}(x_{i})}{F_{S}(k)}$ is strictly less than 1, and $x_{i}<k$, so the CDFs are well-defined.
@@
Since each $X_{i,k}$ has a finite expectation ($\mathbb{E}[X_{i,k}]\leq k<\infty$), Theorem 1 in Imamura and Kato, (2025) implies that $\bm{X}_{k}$ is co-monotonic and therefore VaR additive i.e.:
@@
Theorem 2.2 completes the picture initiated by Imamura and Kato, (2025). For right-sided random variables, VaR sub-additivity at all confidence levels is not merely sufficient but necessary for co-monotonicity, regardless of whether the marginals possess finite expectations. In other words, the only dependence structure compatible with VaR sub-additivity in this setting is the most extreme form of positive dependence, under which diversification provides no benefit of any kind. This result has a clear practical implication: within the class of non-negative risks, seeking sub-additive VaR aggregation is equivalent to accepting that the portfolio components move in perfect lockstep.
@@ 3 VaR super-additivity
Unlike the sub-additivity property of the VaR, the opposite effect – VaR super-additivity – can in fact arise. For instance, consider the case of a counter-monotonic random vector defined as
@@
whose joint CDF is given by the Fréchet lower bound:
@@
For simplicity, take $\alpha=1/2$ and $\theta=1$. The VaRs of the non-identical margins are then given as
${\mathrm{VaR}}_{p}[X]=\frac{p\,(2-p)}{(1-p)^{2}},\quad{\mathrm{VaR}}_{p}\left[\dfrac{1}{X}\right]=\frac{p^{2}}{1-p^{2}}.$
@@
${\mathrm{VaR}}_{p}\left[X+\dfrac{1}{X}\right]=\frac{2\left(1+2p^{2}-p^{4}\right)}{\left(1-p^{2}\right)^{2}}.$
@@
$\frac{2\left(1+2p^{2}-p^{4}\right)}{\left(1-p^{2}\right)^{2}}>\frac{2p\,\left(1+p-p^{2}\right)}{(1-p)^{2}(1+p)},\qquad\forall\,p\in(0,1),$
@@
Although many studies have focused on tail-level VaR super-additivity (Embrechts et al., , 2008, 2009; Zhu et al., , 2023), recent years have witnessed substantial progress on full-range VaR super-additivity, largely analyzed through the lens of first-order stochastic dominance. Formally, this body of work considers random variables $X$ satisfying
$X\leq_{st}\,\sum_{i=1}^{n}\theta_{i}X_{i},$ (3.1)
where $X_{1},\dots,X_{n}$ are independent variables identically distributed to $X$, with $\theta_{i}\in(0,1)$ and $\sum_{i=1}^{n}\theta_{i}=1$.
Within this framework, Ibragimov, (2009) established the stochastic dominance relation for right-sided stable distributions. Subsequently, Chen et al., (2025) relaxed the independence assumption and identified a class of weakly negatively associated super-Pareto risks exhibiting full VaR super-additivity. This class was further extended in Chen and Shneer, (2026), where the analysis relies on the sub-additivity of the function $\xi(x)=-\log F_{X}(1/x)$ i.e. $\xi(x+y)\leq\xi(x)+\xi(y)$, allowing the results to additionally cover risks that are negatively lower orthant dependent. Other contributions, such as Arab et al., (2025) and Müller, (2025), maintain the independence assumption but broaden the classes introduced in Chen and Shneer, (2026) and Ibragimov, (2009). In particular, they introduce the InvSub class – those $X$ for which $\overline{F}_{X}(1/x)$ is sub-additive – and the super-Cauchy class, respectively. Taken together, these results demonstrate that VaR super-additivity is not merely a pathological anomaly but rather arises naturally under economically meaningful and probabilistically coherent conditions.
The simple example presented at the beginning of this section does not satisfy the stochastic dominance formulation in (3.1) and therefore falls outside the classes studied thus far. The objective of this section is to develop a complementary theory to Chen and Shneer, (2026) accommodating non-identical margins and a wider range of dependence structures. Although the class introduced in Arab et al., (2025) contains that of Chen and Shneer, (2026), this inclusion holds only under independence and does not extend to our general setting. Likewise, the super-Cauchy class studied in Müller, (2025) intersects with that of Chen and Shneer, (2026), but it is again restricted to independence and, moreover, includes random variables supported on the entire real line which falls outside our one-sided scope.
Intrinsically, these limitations arise because the stochastic dominance framework in (3.1) does not apply to non-identical risks. To address this gap, we develop a new analytical framework for characterizing VaR super-additivity for right-sided random variables. Our approach allows for heterogeneous marginal distributions and naturally extends the analysis to dependence structures beyond negative lower orthant dependence.
Within this broader setting, nonetheless, VaR super-additivity remains subject to inherent non-integrability requirements. In particular, Theorem 1 of Imamura and Kato, (2025) imposes strong constraints on when super-additivity can happen. We summarize this implication in the following remark.
Remark 3.1.
If all components $X_{i}$ are integrable, i.e. $\mathbb{E}[|X_{i}|]<\infty$ for all $i\in\{1,\dots,n\}$, then Theorem 1 of Imamura and Kato, (2025) implies that VaR super-additivity must degenerate into exact VaR additivity, and $\bm{X}$ must be co-monotonic. This follows directly from their result: if $\bm{X}$ has integrable components and is VaR super-additive, then the reflected variables $\bm{X}$ are integrable and VaR sub-additive, and hence co-monotonic. Since co-monotonicity of $\bm{X}$ is equivalent to co-monotonicity of $\bm{X}$, the conclusion follows.
Remark 3.1 therefore shows that, in order to construct examples of random vectors $\bm{X}$ that are genuinely VaR super-additive, at least one component must be non-integrable. The counter-monotonic example presented earlier illustrates this requirement, as both variables possess infinite means. The converse, however, does not hold. Non-integrability is fundamentally a tail property and, by itself, does not guarantee full VaR super-additivity. The following example illustrates this point.
@@
    we obtain $\mathbb{E}\!\left[\dfrac{1}{1+X}\right]=\dfrac{1}{2}$, while $\mathbb{E}[X]=\infty$ for this Pareto distribution.
@@
    For the sum $S=X+\dfrac{1}{1+X}$, one can show that
@@
    Comparing ${\mathrm{VaR}}_{p}[S]$ with ${\mathrm{VaR}}_{p}[X]+{\mathrm{VaR}}_{p}\!\left[\dfrac{1}{1+X}\right]$ reveals that VaR is super-additive for $p\in\left(0,\dfrac{1}{2}\right],$ and sub-additive for $p\in\left[\dfrac{1}{2},1\right).$
@@
    where $F_{X_{1}}$ and $F_{X_{2}}$ are Pareto (II) marginal distributions with parameters $\alpha_{1}=\alpha_{2}=\theta_{1}=\theta_{2}=1$. The copula $C(u,v)$ is an Ordinal Sum copula (see Example 3.4 in Nelsen, (2010)) given by
    $C(u,v)=\begin{cases}\max\!\left\{u+v-\dfrac{1}{2},\,0\right\},&(u,v)\in\left[0,\dfrac{1}{2}\right]^{2},\\[5.10011pt] \max\!\left\{u+v-1,\,\dfrac{1}{2}\right\},&(u,v)\in\left(\dfrac{1}{2},1\right]^{2},\\[5.10011pt] \min\{u,v\},&\text{otherwise}.\end{cases}$
@@
    Since both margins are Pareto (II) with unit shape then their expectations are infinite. Their common VaR is
@@
    ${\mathrm{VaR}}_{p}[S]=\begin{cases}\dfrac{6+8p^{2}}{\,9-4p^{2}\,},&0<p\leq\dfrac{1}{2},\\[6.80011pt] \dfrac{2-2p\,(1-p)}{p\,(1-p)},&\dfrac{1}{2}<p<1.\end{cases}$
@@
Example 3.2 demonstrates that even in cases where we have (1) counter-monotonic dependence with one non-integrable margin, and (2) two non-counter-monotonic, non-integrable margins, the resulting dependence-margin combination may still exhibit intervals of VaR sub-additivity. Thus, neither a particular dependence structure nor the mere non-integrability of margins is sufficient on its own to guarantee VaR super-additivity.
@@
Our objective, therefore, is to identify a dependence property together with a corresponding marginal behavior that, when combined, imply VaR super-additivity. Such a characterization must be sufficiently general to encompass the class in Chen and Shneer, (2026), yet tractable enough to allow for straightforward verification.
@@
is SD, then $\bm{X}$ is VaR super-additive.
@@
When $\bm{X}$ is negative lower orthant dependent (NLOD) (Block et al., , 1982; Joe, , 1997), that is
@@
Remark 3.9.
If we define a random vector $\bm{Y}=(Y_{1},\dots,Y_{n})$ via $Y_{i}=\theta_{i}X_{i}$, where $\theta_{i}\in(0,1)$, $\sum_{i=1}^{n}\theta_{i}=1$, and $X_{1},\dots,X_{n}$ are identically distributed NLOD variables (consequently, $\bm{Y}$ is NLOD). VaR super-additivity of $\bm{Y}$ is therefore equivalent to the stochastic dominance relation in (3.1).
Setting $y_{i}=\theta_{i}s$ with $s=\sum_{i=1}^{n}y_{i},\,\theta_{i}=y_{i}/s$, the SD property of the function $\Phi$ in Equation (3.2) for $\bm{Y}$ becomes equivalent to
$F_{X}(s)\geq\prod_{i=1}^{n}F_{X}\left(\frac{s}{\theta_{i}}\right),\quad\forall s\in[0,\infty),\,\mathrm{and}\,\,\theta_{i}\in(0,1)\,\,\mathrm{with}\,\sum_{i=1}^{n}\theta_{i}=1,$
where $X$ is a random variable identically distributed to $X_{i}$.
This condition is exactly the sub-additivity property of the function $\xi(x)=-\log F_{X}(1/x)$. Consequently, together with the fact that NLOD is a special case of NSD as established in Proposition 3.6, it follows that $X$ corresponds to the class studied in Chen and Shneer, (2026).
The NSD property captures the dependence requirement on the joint distribution of $\bm{X}$ that ensures VaR super-additivity. We note, in passing, that the dependence structure used in part (2) of Example 3.2 is weaker than NSD, specifically it fails the NSD property at $t\in\left(\dfrac{7}{10},1+\sqrt{2}\right)$. This contributed, though was not strictly required, to the failure of VaR super-additivity in that example. Nonetheless, by definition, NSD is a relatively weak form of negative dependence and is strictly implied by NLOD.
@@
${\mathrm{VaR}}_{p}[X]={\mathrm{VaR}}_{p}\left[\dfrac{1}{X}\right]=\frac{p}{1-p},$
@@
We collect below several familiar continuous distributions that are widely used to model claim severities in insurance whose associated functions $\phi_{i}$ are non-increasing.
@@
    $F_{X_{i}}(x_{i})=\exp\left(-\left(\dfrac{x_{i}}{\theta_{i}}\right)^{-\alpha_{i}}\right),\qquad x_{i}\geq 0,\ \alpha_{i},\theta_{i}>0,$
@@
    $F_{X_{i}}(x_{i})=1-\left(\dfrac{\theta_{i}}{\theta_{i}+x_{i}}\right)^{\alpha_{i}},\qquad x_{i}\geq 0,\ \alpha_{i},\theta_{i}>0,$
@@
    $\phi_{i}(x_{i})=x_{i}\log\!\left(1-\left(\dfrac{\theta_{i}}{\theta_{i}+x_{i}}\right)^{\alpha_{i}}\right).$
@@
    $\phi_{i}^{{}^{\prime}}(x_{i})$ $=\log F_{X_{i}}(x_{i})+\dfrac{x_{i}f_{X_{i}}(x_{i})}{F_{X_{i}}(x_{i})}$
    $=\log F_{X_{i}}(x_{i})+\dfrac{\alpha_{i}x_{i}\overline{F}_{X_{i}}(x_{i})}{(\theta_{i}+x_{i})F_{X_{i}}(x_{i})}.$
@@
    $\lim_{x_{i}\to\infty}\dfrac{\phi_{i}^{{}^{\prime}}(x_{i})}{\overline{F}_{X_{i}}(x_{i})}=\alpha_{i}-1.$
@@
    $\phi_{i}^{{}^{\prime}}(x_{i})$ $=-\int_{F_{X_{i}}(x_{i})}^{1}\dfrac{1}{w}\,\mathrm{d}w+\dfrac{\alpha_{i}x_{i}}{(\theta_{i}+x_{i})F_{X_{i}}(x_{i})}\int_{F_{X_{i}}(x_{i})}^{1}\mathrm{d}w$
    $=-\int_{F_{X_{i}}(x_{i})}^{1}\dfrac{(\theta_{i}+x_{i})F_{X_{i}}(x_{i})-\alpha_{i}x_{i}w}{w(\theta_{i}+x_{i})F_{X_{i}}(x_{i})}\,\mathrm{d}w.$
@@
    Applying the mean value theorem to $t\mapsto t^{\alpha_{i}}$ on $\left[\dfrac{\theta_{i}}{\theta_{i}+x_{i}},1\right]$ yields the inequality and thus the desired non-positivity.
@@
    $\phi_{i}^{\prime}(x_{i})=\log\!\left(\mathrm{erfc}\!\left(\sqrt{\frac{\theta_{i}}{2x_{i}}}\right)\right)+\frac{\sqrt{\dfrac{\theta_{i}}{2\pi x_{i}}}\exp\!\left(-\dfrac{\theta_{i}}{2x_{i}}\right)}{\mathrm{erfc}\!\left(\sqrt{\dfrac{\theta_{i}}{2x_{i}}}\right)}.$
    Let $t=\sqrt{\dfrac{\theta_{i}}{2x_{i}}}$. Then $\phi_{i}^{\prime}(x_{i})\leq 0$ is equivalent to $\psi_{i}(t)\leq 0$, where
@@
    the classical Mills ratio bound (Mills, , 1926). Thus $\phi_{i}^{\prime}(x_{i})\leq 0$ for all $x_{i}\geq 0$ i.e. $\phi_{i}$ is non-increasing on $0,\infty)$.
@@
    $F_{X_{i}}(x_{i})=\left(\dfrac{x_{i}}{1+x_{i}}\right)^{\alpha_{i}},\qquad x_{i}\geq 0,\ \alpha_{i}>0,$
@@
    $\phi_{i}(x_{i})=\alpha_{i}x_{i}\log\!\left(\dfrac{x_{i}}{1+x_{i}}\right).$